OpenStudy (anonymous):

A line with an inclination passes through (-5/2, -9/2), what is the x-coordinate of a point on the line if its corresponding y-coordinate is 6?

OpenStudy (anonymous):

@Michele_Laino can you help me? :)

OpenStudy (michele_laino):

if we write the line3 as below: \[\Large ax + by + c = 0\] we get, by substitution: \[\Large 2c = 5a + 9b\] next I substitute \(y=6\) into the first equation: \[\Large a{x_0} + 6b + c = 0\] then I substitute \(c\): \[\Large a\left( {{x_0} + 5} \right) + 15b = 0\]

OpenStudy (anonymous):

why did you put (Xo + 5 )?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

sorry I have made an error. If I substitute the value for \(c\), I get: \[\Large \begin{gathered} a{x_0} + 6b + \frac{{5a + 9b}}{2} = 0 \hfill \\ \hfill \\ 2a{x_0} + 12b + 5a + 9b = 0 \hfill \\ \hfill \\ a\left( {2{x_0} + 5} \right) + 21b = 0 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

where \(x_0\) is the requested x-coordinates

OpenStudy (michele_laino):

x-coordinate*

OpenStudy (anonymous):

then what should i do? i'm sorry :(

OpenStudy (michele_laino):

I also don't know I think that we need to know another quantity, for example, the slope of the line

OpenStudy (michele_laino):

for example, do the line pass at the origin \((0,0)\)

OpenStudy (michele_laino):

does*